# Hyperelliptic four-manifolds defined by vector-colorings of simple polytopes

### v2

@@ 4.7. Geometric hyperelliptic manifolds $N(P,\Lambda)$
Construction 4.45 is a direct generalization to dimensions $n>3$ of the construction of geometric hyperelliptic $3$-manifolds in [M90, VM99M, VM99S2] based on Hamiltonian cycles, theta- and $K_{4}$-subgraphs in right-angled 33-polytopes.
@@
Theorem 4.39 implies that for $n\leqslant 4$ and any linearly independent vector-coloring $\Lambda$ of a right-angled $n$-polytope $P$ if $N(P,\Lambda)$ admits a hyperelliptic involution in $\mathbb{Z}_{2}^{r}$, then this involution, the coloring $\Lambda$ and $N(P,\Lambda)$ can be obtained by Construction 4.45.
In this section we will discuss which geometries $\mathbb{X}$ admit hyperelliptic manifolds obtained by Construction 4.45.
@@
In particular, each vertex of $P$ either is a vertex of $C$ or lies on its $1$-face. Thus, $C$ should have $1$-faces. This is possible only for $k\in\{n-1,n,n+1\}$. Moreover, $\mathcal{C}^{1}(n,n-1)$ is a circle, $\mathcal{C}^{1}(n,n-1)$ consists of two vertices and $n$ multiple edges, and $\mathcal{C}^{1}(n,n-1)$ is the complete graph on $n$ vertices (see [E24, Example 8.17]). If a vertex $v$ of $P$ is a vertex of $C$, then it is the intersection of $n$ different facets of $C$, and it does not lie in defining faces. If $v$ lies in a $1$-skeleton of $C$, then two facets of $P$ containing $v$ lie in the same facet of $C$. Then their intersection is a defining face.
For k=n1>1k=n-1>1 we have $m=f_{n-1}(P)={r\choose 1}+\sum\limits_{q=1}^{s}f_{n-2}(M_{q})=r-s$. Thus, s=mrs=m-r is the total number of defining faces. ∎
@@
The $n$-cube $I^{n}$ does not admit Hamiltonian $\mathcal{C}(n,k)$-subcomplexes for $n>3$. In particular, there are no locally Euclidean hyperelliptic $n$-manifolds $N(P,\Lambda)$, $n>3$, obtained by Construction 4.45.
@@
The cube has $2^{n}$ vertices and $2n$ facets. Any (n2)(n-2)-face is $I^{n-2}$ and has $2^{n-2}$ vertices. If $I^{n}$ has a Hamiltonian $\mathcal{C}(n,k)$-subcomplex, then $2^{n}=V_{k}+(2n-k)2^{n-2}$, where $n-1\leqslant k\leqslant n+1$. We have $2n-k\geqslant n-1$. Thus, the right part is at least $V_{k}+(n-1)2^{n-2}$. If $n\geqslant 5$, then it is at least $V_{k}+4\cdot 2^{n-2}=V_{k}+2^{n}\geqslant 2^{n}$, where the equality holds only if $n=5$ and $k=n+1$. But in this case $V_{k}=n+1=6$ and we obtain a contradiction. If $n=4$, then we have $16=V_{k}+(8-k)\cdot 4$. For $k=3$ we have $16=0+5\cdot 4=20$, for $k=4$ we have $16=2+4\cdot 4=18$, and for $k=5$ we have $16=5+3\cdot 4=17$. A contradiction. ∎
@@
Let and $C$ be a Hamiltonian $\mathcal{C}(n,n-1)$-subcomplex in $\partial P$. Then any its defining face is an (n2)(n-2)-face of $P$ admitting a coloring in (n2)(n-2) colors such that adjacent facets have different colors.
@@
The first example of a small cover $N(P,\Lambda)$ over a $4$-polytope $P$ with a hyperelliptic involution in $\mathbb{Z}_{2}^{4}$ was build by Alexei Koretskii, see [K24].
@@
If a $4$-polytope $P$ admits a Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then $P$ has at least one triangular or quadrangular 22-face. In particular, there are no hyperbolic ($\mathbb{X}=\mathbb{L}^{4}$) hyperelliptic $4$-manifolds $N(P,\Lambda)$ obtained by Construction 4.45. Moreover, for all $n\geqslant 4$ there are no hyperbolic ($\mathbb{X}=\mathbb{L}^{n}$) $n$-manifolds $N(P,\Lambda)$ obtained by Construction 4.45.
@@
Let $p_{k}$ be the number of $k$-gonal 22-faces of $P$, and $f_{i}$ be the total number of its $i$-faces (in particular, $f_{3}=m$). We have the Euler formula $f_{0}-f_{1}+f_{2}-f_{3}=0$. Also $4f_{0}=2f_{1}$ since $P$ is simple. Also $\sum\limits_{k}kp_{k}={4\choose 2}f_{0}=6f_{0}$. Then $f_{2}=f_{1}-f_{0}+f_{3}=f_{0}+f_{3}=f_{0}+m$, and
@@
If a $4$-polytope $P=Q\times I$ admits a Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then at least one of the defining $2$-faces of $P$ is a triangle. In particular, there are no hyperelliptic $4$-manifolds $N(P,\Lambda)$ with geometries $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$, and $\mathbb{R}^{4}$ obtained by Construction 4.45.
@@
For $n\geqslant 4$ the simplex $\Delta^{n}$ up to symmetries has a unique proper Hamiltonian $\mathcal{C}(n,r)$-subcomplex defined by a single (n2)(n-2)-face $\Delta^{n-2}$. For this subcomplex $r=n$. It corresponds to a hyperelliptic involution on the manifold $\mathbb{R}\mathcal{Z}_{\Delta^{n}}\simeq S^{n}$. In particular, the geometry $\mathbb{S}^{n}$ arises in Construction 4.45. The branch set of the covering $S^{n}\to S^{n}$ is the sphere $S^{n-2}$.
@@
For $n\geqslant 4$ the prism $\Delta^{n-1}\times I$ up to symmetries admits exactly $3$ Hamiltonian $\mathcal{C}(n,r)$-subcomplexes: a unique $\mathcal{C}(n,n+1)$-subcomplex and two $\mathcal{C}(n,n)$-subcomplexes. In particular, the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ arises in Construction 4.45. For the $\mathcal{C}(n,n+1)$-subcomplex and the $\mathcal{C}(n,n)$-subcomplexes the branch sets of the coverings $S^{n-1}\times S^{1}\to S^{n}$ and $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ are disjoint unions of two spheres $S^{n-2}$.
@@
Construction 4.62 (Cutting off a face).
Let $P=\{\boldsymbol{a}_{i}\boldsymbol{x}+b_{i}\geqslant 0,i=1,\dots,m\}\subset\mathbb{R}^{n}$ be a simple $n$-polytope and $G=F_{i_{1}}\cap\dots\cap F_{i_{k}}$ be its (nk)(n-k)-face, $k>1$. Set $\boldsymbol{a}_{G}=\sum\limits_{s=1}^{k}\boldsymbol{a}_{i_{s}}$ and $b_{G}=\sum\limits_{s=1}^{k}b_{i_{s}}$. Take the halfspace $\mathcal{H}_{\geqslant\varepsilon}=\{\boldsymbol{a}_{G}\boldsymbol{x}+b_{G}\geqslant\varepsilon\}$. For small $\varepsilon\geqslant 0$ the intersection $P_{G,\varepsilon}=P\cap\mathcal{H}_{\geqslant\varepsilon}$ is a simple $n$-polytope. It has an additional facet $F$ corresponding to the intersection of $P$ with the hyperplane $\mathcal{H}_{\varepsilon}=\partial\mathcal{H}_{\geqslant\varepsilon}$. We have $F\cap F_{j_{1}}\cap\dots\cap F_{j_{l}}\neq\varnothing$ if and only if $F_{j_{1}}\cap\dots\cap F_{j_{l}}$ is a face of $P$ intersecting $G$ and not lying in $G$. That is, $F_{j_{1}}\cap\dots\cap F_{j_{l}}\cap F_{i_{1}}\cap\dots\cap F_{i_{k}}\neq\varnothing$ and $\{i_{1},\dots,i_{k}\}\not\subset\{j_{1},\dots,j_{l}\}$.
@@
Any (n2)(n-2)-face $F\cap F_{i_{q}}\simeq G\times\Delta^{k-2}$ of $P_{G,\varepsilon}$ defines a Hamiltonian subcomplex $C\simeq\partial P$.
@@
We have $\mathcal{H}(0)=\mathcal{H}_{\varepsilon}={\rm aff}(F)$ and $\mathcal{H}(\frac{\pi}{2})={\rm aff}(F_{i_{q}})$. Moreover, for small $\delta>0$ at each moment $\delta\leqslant\varphi<\frac{\pi}{2}$ the polytope $P\cap\mathcal{H}_{\geqslant 0}(\varphi)$ has the same combinatorial type $P_{\varepsilon}$ and for any $\delta\leqslant\varphi_{1}<\varphi_{2}\leqslant\frac{\pi}{2}$ the polytope $P\cap\mathcal{H}_{\leqslant 0}(\varphi_{1})\cap\mathcal{H}_{\geqslant}(\varphi_{2})$ has the same combinatorial type $G\times\Delta^{k}\simeq P\cap\mathcal{H}_{\leqslant\varepsilon}$. In particular, the polytope $P\cap\mathcal{H}_{\leqslant 0}(-\delta)\cap\mathcal{H}_{\geqslant 0}(0)$ is combinatorially equivalent to $P\cap\mathcal{H}_{\leqslant 0}(-\delta)\cap\mathcal{H}_{\geqslant}(\frac{\pi}{2})$. Then the piecewise linear homeomorphism defined using the barycentric subdivisions preserves the face structures and is identical on $P\cap\mathcal{H}(-\delta)$. This homeomorphism together with the identical mapping on $P\cap\mathcal{H}_{\geqslant 0}(-\delta)$ defines the desired equivalence $C\simeq\partial P$. ∎
@@
For any $p,q\geqslant 1$ the face $\Delta^{p-1}\times\Delta^{q-1}$ defines a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex in $P^{n}=\Delta^{p}\times\Delta^{q}$. In particular, for $p,q\geqslant 2$ the geometry $\mathbb{S}^{p}\times\mathbb{S}^{q}$ arises in Construction 4.45. (For $p=1$ or $q=1$ we obtain the polytope $\Delta^{n-1}\times I$ and the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ already covered by Proposition 4.61.) For $p,q\geqslant 2$ the branch set of the coverings $S^{p}\times S^{q}\to S^{n}$ is homeomorphic to $S^{p-1}\times S^{q-1}$.
@@
If an $n$-polytope $P$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect a facet $F_{i}\simeq\Delta^{n-1}$, then for any $k\geqslant 1$ the polytope $\Delta^{k}\times P$ has a Hamiltonian $\mathcal{C}(n+k,n+k+1)$-subcomplex with defining faces $\Delta^{k-1}\times F_{i}$ and $\Delta^{k}\times M_{1}$, $\dots$, $\Delta^{k}\times M_{q}$.
@@
Indeed, the faces $\Delta^{k}\times M_{1}$, $\dots$, $\Delta^{k}\times M_{q}$ define the subcomplex equivalent to $\partial(\Delta^{k}\times\Delta^{n})$, and the face $\Delta^{k-1}\times F_{i}\simeq\Delta^{k-1}\times\Delta^{n-1}$ transforms it to $\partial\Delta^{n+k}$. ∎
@@
For any $p,q\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times\Delta^{q}\times I$ has a Hamiltonian $C(n,n+1)$-subcomplex defined by two faces $\Delta^{p-1}\times\Delta^{q}\times\{0\}$ and $\Delta^{p}\times\Delta^{q-1}\times\{1\}$. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{q}\times\mathbb{R}$, $p,q\geqslant 2$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$, $p\geqslant 2$, and $\mathbb{R}^{3}$ arise in Construction 4.45. The branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ is $S^{p-1}\times S^{q}\sqcup S^{p}\times S^{q-1}$.
@@
Let $\Delta^{2}={\rm conv}\{v_{1},v_{2},v_{3}\}$. Then the faces ${\rm conv}\{v_{1},v_{2}\}\times\Delta^{n-3}$ and $v_{3}\times\Delta^{n-2}$ define a Hamiltonian $\mathcal{C}(n,n)$-subcomplex $C$ in $\Delta^{2}\times\Delta^{n-2}$ for $n\geqslant 3$. The branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ is a disjoint union of $S^{n-2}$ and $S^{1}\times S^{n-3}$.
@@
Let $P_{k}$ be a $k$-gon, $k\geqslant 3$. Then for $p\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times P_{k}$ admits Hamiltonian $\mathcal{C}(n,n)$- and $\mathcal{C}(n,n+1)$-subcomplexes $C_{n}$ and $C_{n+1}$. In particular, taking a right-angled triangle in $\mathbb{S}^{2}$, a square in $\mathbb{R}^{2}$ and a right-angled $k$-gon, $k\geqslant 5$, in $\mathbb{L}^{2}$ we see that the geometries $\mathbb{S}^{p}\times\mathbb{S}^{2}$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$ and $\mathbb{S}^{p}\times\mathbb{L}^{2}$, $p\geqslant 2$, arise in Construction 4.45. The branch set of the covering $N(P,\widetilde{\Lambda}_{C_{n+1}})\to S^{n}$ is a disjoint union of 2(k3)2(k-3) copies of $S^{n-2}$ and one $S^{n-3}\times S^{1}$, and for the covering $N(P,\widetilde{\Lambda}_{C_{n}})\to S^{n}$ it is a disjoint union of (k2)(k-2)-copies of $S^{n-2}$ and one $S^{n-3}\times S^{1}$.
@@
On Fig. 1 we show a set of five disjoint quadrangles defining a Hamiltonian $\mathcal{C}(4,5)$-subcomplex in $P_{5}\times P_{5}$. In particular, the geometry $\mathbb{L}^{2}\times\mathbb{L}^{2}$ arises in Construction 4.45. It follows from Lemma 4.60 that the branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{4}$ is a disjoint union of five tori $\mathbb{T}^{2}=S^{1}\times S^{1}$.

### v3

@@ 4.7. Geometric hyperelliptic manifolds $N(P,\Lambda)$
Construction 4.46 is a direct generalization to dimensions $n>3$ of the construction of geometric hyperelliptic $3$-manifolds in [M90, VM99M, VM99S2] based on Hamiltonian cycles, theta- and $K_{4}$-subgraphs in 11-skeletons of right-angled 33-polytopes.
@@
Theorem 4.40 implies that for $n\leqslant 4$ and any linearly independent vector-coloring $\Lambda$ of a right-angled $n$-polytope $P$ if $N(P,\Lambda)$ admits a hyperelliptic involution in $\mathbb{Z}_{2}^{r}$, then this involution, the coloring $\Lambda$ and $N(P,\Lambda)$ can be obtained by Construction 4.46.
In this section we will discuss which geometries $\mathbb{X}$ admit hyperelliptic manifolds obtained by Construction 4.46.
@@
In particular, each vertex of $P$ either is a vertex of $C$ or lies on its $1$-face. Thus, $C$ should have $1$-faces. This is possible only for $k\in\{n-1,n,n+1\}$. Moreover, $\mathcal{C}^{1}(n,n-1)$ is a circle, $\mathcal{C}^{1}(n,n-1)$ consists of two vertices and $n$ multiple edges, and $\mathcal{C}^{1}(n,n+1)$ is the complete graph on $n$ vertices (see [E24, Example 8.17]). If a vertex $v$ of $P$ is a vertex of $C$, then it is the intersection of $n$ different facets of $C$, and it does not lie in defining faces. If $v$ lies inside a $1$-face of $C$, then two facets of $P$ containing $v$ lie in the same facet of $C$. Then their intersection is a defining face.
For k=n1>1k=n-1>1 we have $m=f_{n-1}(P)={r\choose 1}+\sum\limits_{q=1}^{s}f_{n-2}(M_{q})=r+s$. Thus, s=mrs=m-r is the total number of defining faces. ∎
@@
The $n$-cube $I^{n}$ does not admit Hamiltonian $\mathcal{C}(n,k)$-subcomplexes for $n>3$. In particular, there are no locally Euclidean hyperelliptic $n$-manifolds $N(P,\Lambda)$, $n>3$, obtained by Construction 4.46.
@@
The cube has $2^{n}$ vertices and $2n$ facets. Any (n2)(n-2)-face is $I^{n-2}$ and has $2^{n-2}$ vertices. If $I^{n}$ has a Hamiltonian $\mathcal{C}(n,k)$-subcomplex, then $2^{n}=V_{n,k}+(2n-k)2^{n-2}$, where $n-1\leqslant k\leqslant n+1$. We have $2n-k\geqslant n-1$. Thus, the right part is at least $V_{n,k}+(n-1)2^{n-2}$. If $n\geqslant 5$, then it is at least $V_{n,k}+4\cdot 2^{n-2}=V_{n,k}+2^{n}\geqslant 2^{n}$, where the equality holds only if $n=5$ and $k=n+1$. But in this case $V_{n,k}=n+1=6$ and we obtain a contradiction. If $n=4$, then we have $16=V_{n,k}+(8-k)\cdot 4$. For $k=3$ we have $16=0+5\cdot 4=20$, for $k=4$ we have $16=2+4\cdot 4=18$, and for $k=5$ we have $16=5+3\cdot 4=17$. A contradiction. ∎
@@
Let and $C$ be a Hamiltonian $\mathcal{C}(n,n-1)$-subcomplex in $\partial P$. Then any its defining face is an (n2)(n-2)-polytope admitting a coloring of facets in (n2)(n-2) colors such that adjacent facets have different colors.
@@
The first example of a small cover $N(P,\Lambda)$ over a $4$-polytope $P$ with a hyperelliptic involution in $\mathbb{Z}_{2}^{4}$ was build by Alexei Koretskii, see [K24B] and [K24].
@@
If a $4$-polytope $P$ admits a Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then $P$ has at least one triangular or quadrangular 22-face. In particular, there are no hyperbolic ($\mathbb{X}=\mathbb{L}^{4}$) hyperelliptic $4$-manifolds $N(P,\Lambda)$ obtained by Construction 4.46. Moreover, for all $n\geqslant 4$ there are no hyperbolic ($\mathbb{X}=\mathbb{L}^{n}$) $n$-manifolds $N(P,\Lambda)$ obtained by Construction 4.46.
@@
Let $p_{k}$ be the number of $k$-gonal 22-faces of $P$, and $f_{i}$ be the total number of its $i$-faces (in particular, $f_{3}=m$). We have the Euler-Poincare formula $f_{0}-f_{1}+f_{2}-f_{3}=0$. Also $4f_{0}=2f_{1}$ since $P$ is simple. Also $\sum\limits_{k}kp_{k}={4\choose 2}f_{0}=6f_{0}$. Then $f_{2}=f_{1}-f_{0}+f_{3}=f_{0}+f_{3}=f_{0}+m$, and
@@
If a $4$-polytope $P=Q\times I$ admits a Hamiltonian $\mathcal{C}(4,r)$-subcomplex, then at least one of the defining $2$-faces of $P$ is a triangle. In particular, there are no hyperelliptic $4$-manifolds $N(P,\Lambda)$ with geometries $\mathbb{L}^{3}\times\mathbb{R}$, $\mathbb{L}^{2}\times\mathbb{R}^{2}$, and $\mathbb{R}^{4}$ obtained by Construction 4.46.
@@
For $n\geqslant 4$ the simplex $\Delta^{n}$ up to symmetries has a unique proper Hamiltonian $\mathcal{C}(n,r)$-subcomplex defined by a single (n2)(n-2)-face $\Delta^{n-2}$. For this subcomplex $r=n$. It corresponds to a hyperelliptic involution on the manifold $\mathbb{R}\mathcal{Z}_{\Delta^{n}}\simeq S^{n}$. In particular, the geometry $\mathbb{S}^{n}$ arises in Construction 4.46. The branch set of the covering $S^{n}\to S^{n}$ is the sphere $S^{n-2}$.
@@
For $n\geqslant 4$ the prism $\Delta^{n-1}\times I$ up to symmetries admits exactly $3$ Hamiltonian $\mathcal{C}(n,r)$-subcomplexes: a unique $\mathcal{C}(n,n+1)$-subcomplex and two $\mathcal{C}(n,n)$-subcomplexes. In particular, the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ arises in Construction 4.46. For the $\mathcal{C}(n,n+1)$-subcomplex and the $\mathcal{C}(n,n)$-subcomplexes the branch sets of the coverings $S^{n-1}\times S^{1}\to S^{n}$ and $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ are disjoint unions of two spheres $S^{n-2}$.
@@
Construction 4.63 ((Cutting off a face)).
Let $P=\{\boldsymbol{a}_{i}\boldsymbol{x}+b_{i}\geqslant 0,i=1,\dots,m\}\subset\mathbb{R}^{n}$ be a simple $n$-polytope and $G=F_{i_{1}}\cap\dots\cap F_{i_{k}}$ be its (nk)(n-k)-face, $k>1$. Set $\boldsymbol{a}_{G}=\sum\limits_{s=1}^{k}\boldsymbol{a}_{i_{s}}$ and $b_{G}=\sum\limits_{s=1}^{k}b_{i_{s}}$. Take the halfspace $\mathcal{H}_{\geqslant\varepsilon}=\{\boldsymbol{a}_{G}\boldsymbol{x}+b_{G}\geqslant\varepsilon\}$. For small $\varepsilon>0$ the intersection $P_{G,\varepsilon}=P\cap\mathcal{H}_{\geqslant\varepsilon}$ is a simple $n$-polytope. It has an additional facet $F$ corresponding to the intersection of $P$ with the hyperplane $\mathcal{H}_{\varepsilon}=\partial\mathcal{H}_{\geqslant\varepsilon}$. We have $F\cap F_{j_{1}}\cap\dots\cap F_{j_{l}}\neq\varnothing$ if and only if $F_{j_{1}}\cap\dots\cap F_{j_{l}}$ is a face of $P$ intersecting $G$ and not lying in $G$. That is, $F_{j_{1}}\cap\dots\cap F_{j_{l}}\cap F_{i_{1}}\cap\dots\cap F_{i_{k}}\neq\varnothing$ and $\{i_{1},\dots,i_{k}\}\not\subset\{j_{1},\dots,j_{l}\}$.
@@
Any (n2)(n-2)-face $F\cap F_{i_{q}}\simeq G\times\Delta^{k-2}$, $i_{q}\in\{i_{1},\dots,i_{k}\}$, of $P_{G,\varepsilon}$ defines a Hamiltonian subcomplex $C\simeq\partial P$.
@@
We have $\mathcal{H}(0)=\mathcal{H}_{\varepsilon}={\rm aff}(F)$ and $\mathcal{H}(\frac{\pi}{2})={\rm aff}(F_{i_{q}})$. Moreover, for small $\delta>0$ at each moment $\delta\leqslant\varphi<\frac{\pi}{2}$ the polytope $P\cap\mathcal{H}_{\geqslant 0}(\varphi)$ has the same combinatorial type $P_{\varepsilon}$ and for any $\delta\leqslant\varphi_{1}<\varphi_{2}\leqslant\frac{\pi}{2}$ the polytope $P\cap\mathcal{H}_{\leqslant 0}(\varphi_{1})\cap\mathcal{H}_{\geqslant 0}(\varphi_{2})$ has the same combinatorial type $G\times\Delta^{k}\simeq P\cap\mathcal{H}_{\leqslant\varepsilon}$. In particular, the polytope $P\cap\mathcal{H}_{\leqslant 0}(-\delta)\cap\mathcal{H}_{\geqslant 0}(0)$ is combinatorially equivalent to $P\cap\mathcal{H}_{\leqslant 0}(-\delta)\cap\mathcal{H}_{\geqslant 0}(\frac{\pi}{2})$. Then the piecewise linear homeomorphism defined using the barycentric subdivisions preserves the face structures and is identical on $P\cap\mathcal{H}(-\delta)$. This homeomorphism together with the identical mapping on $P\cap\mathcal{H}_{\geqslant 0}(-\delta)$ defines the desired equivalence $C\simeq\partial P$. ∎
@@
For any $p,q\geqslant 1$ the face $\Delta^{p-1}\times\Delta^{q-1}$ defines a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex in $P^{n}=\Delta^{p}\times\Delta^{q}$. In particular, for $p,q\geqslant 2$ the geometry $\mathbb{S}^{p}\times\mathbb{S}^{q}$ arises in Construction 4.46. (For $p=1$ or $q=1$ we obtain the polytope $\Delta^{n-1}\times I$ and the geometry $\mathbb{S}^{n-1}\times\mathbb{R}$ already covered by Proposition 4.62.) For $p,q\geqslant 2$ the branch set of the coverings $S^{p}\times S^{q}\to S^{n}$ is homeomorphic to $S^{p-1}\times S^{q-1}$.
@@
If an $n$-polytope $P$ has a Hamiltonian $\mathcal{C}(n,n+1)$-subcomplex such that its defining faces $\{M_{1},\dots,M_{s}\}$ do not intersect a facet $F_{i}\simeq\Delta^{n-1}$, then for any $k\geqslant 1$ the polytope $\Delta^{k}\times P$ has a Hamiltonian $\mathcal{C}(n+k,n+k+1)$-subcomplex with defining faces $\Delta^{k-1}\times F_{i}$ and $\Delta^{k}\times M_{1}$, $\dots$, $\Delta^{k}\times M_{s}$.
@@
Indeed, the faces $\Delta^{k}\times M_{1}$, $\dots$, $\Delta^{k}\times M_{s}$ define the subcomplex equivalent to $\partial(\Delta^{k}\times\Delta^{n})$, and the face $\Delta^{k-1}\times F_{i}\simeq\Delta^{k-1}\times\Delta^{n-1}$ transforms it to $\partial\Delta^{n+k}$. ∎
@@
For any $p,q\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times\Delta^{q}\times I$ has a Hamiltonian $C(n,n+1)$-subcomplex defined by two faces $\Delta^{p-1}\times\Delta^{q}\times\{0\}$ and $\Delta^{p}\times\Delta^{q-1}\times\{1\}$. In particular, the geometries $\mathbb{S}^{p}\times\mathbb{S}^{q}\times\mathbb{R}$, $p,q\geqslant 2$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$, $p\geqslant 2$, and $\mathbb{R}^{3}$ arise in Construction 4.46. The branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ is $S^{p-1}\times S^{q}\sqcup S^{p}\times S^{q-1}$.
@@
Let $\Delta^{2}={\rm conv}\{v_{1},v_{2},v_{3}\}$. Then the faces ${\rm conv}\{v_{1},v_{2}\}\times\Delta^{n-3}$ and $v_{3}\times\Delta^{n-2}$ define a Hamiltonian $\mathcal{C}(n,n)$-subcomplex $C$ in $\Delta^{2}\times\Delta^{n-2}$ for $n\geqslant 3$. The branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{n}$ is a disjoint union of $S^{1}\times S^{n-3}$ and $S^{n-2}$ .
@@
Let $P_{k}$ be a $k$-gon, $k\geqslant 3$. Then for $p\geqslant 1$ the polytope $P^{n}=\Delta^{p}\times P_{k}$ admits Hamiltonian $\mathcal{C}(n,n)$- and $\mathcal{C}(n,n+1)$-subcomplexes $C_{n}$ and $C_{n+1}$. In particular, taking a right-angled triangle in $\mathbb{S}^{2}$, a square in $\mathbb{R}^{2}$ and a right-angled $k$-gon, $k\geqslant 5$, in $\mathbb{L}^{2}$ we see that the geometries $\mathbb{S}^{p}\times\mathbb{S}^{2}$, $\mathbb{S}^{p}\times\mathbb{R}^{2}$ and $\mathbb{S}^{p}\times\mathbb{L}^{2}$, $p\geqslant 2$, arise in Construction 4.46. The branch set of the covering $N(P,\widetilde{\Lambda}_{C_{n+1}})\to S^{n}$ is a disjoint union of 2(k3)2(k-3) copies of $S^{n-2}$ and one $S^{n-3}\times S^{1}$, and for the covering $N(P,\widetilde{\Lambda}_{C_{n}})\to S^{n}$ it is a disjoint union of (k2)(k-2) copies of $S^{n-2}$ and one $S^{n-3}\times S^{1}$.
@@
On Fig. 1 we show a set of five disjoint quadrangles defining a Hamiltonian $\mathcal{C}(4,5)$-subcomplex in $P_{5}\times P_{5}$. In particular, the geometry $\mathbb{L}^{2}\times\mathbb{L}^{2}$ arises in Construction 4.46. It follows from Lemma 4.61 that the branch set of the covering $N(P,\widetilde{\Lambda}_{C})\to S^{4}$ is a disjoint union of five tori $\mathbb{T}^{2}=S^{1}\times S^{1}$. Comparing Fig. 1 and [M24, Fig. 13 and 17] it can be proved that these five tori coincide with the Ivanšić link [I04, I12], and the manifold $N(P,\widetilde{\Lambda}_{C})$ coincides with its double branched covering space $W$ from [M24, Section 2.2], where on this manifolds the geometric structure $\mathbb{L}^{2}\times\mathbb{L}^{2}$ is also defined. The complement to the Ivanšić link in $S^{4}$ has a complete hyperbolic structure obtained by glueing $32$ copies of the right-angled hyperbolic $4$-polytope of finite volume $P^{4}\subset\mathbb{L}^{4}$. It has $5$ ideal and $5$ finite vertices [M24, Section 1.2]. The dual polytope coincides with the hypersimplex $\Delta(5,2)$. Each facet of $P$ is a triangular bipyramid and has three ideal and two finite vertices. The Ivanšić link is a $4$-dimensional generalisation of the Borromean rings: each two tori form a trivial link, but each three do not [M24, Theorem 1]. The author is grateful to Vladimir Gorchakov for pointing out a connection between this example and the work [M24].
@@
Remark 4.77.
In [E25] there is a construction of a $3$-dimensional generalisation of the Borromean rings: each Hamiltonian theta-subgraph $\Gamma$ in the $1$-skeleton of a compact right-angled hyperbolic $3$-polytope $P$ such that each edge of $P$ not lying in $\Gamma$ connects vertices on different paths of $\Gamma$ corresponds to a link $L$ of circles in $S^{3}$ with the complement having a complete hyperbolic structure. Any two circles of $L$ form a trivial link, but $L$ contains the Borromean rings and therefore is nontrivial.